OpenStudy (issy14):

A wheel of diameter of .35m spins with an angular velocity of 188 radians per second. Find the linear velocity of the wheel.

OpenStudy (issy14):

No idea, on how to solve this. I know the formula for liner velocity is v=dt. How do i get the time from the angular velocity?

OpenStudy (anonymous):

D two parts. 188 radians per sec is (188/2 pi) revolutions per sec = 29.9/sec each revolution gives distance of pi D, the circumference so, v = pi D (29.9/sec) = 32.9 m/s

OpenStudy (issy14):

@douglaswinslowcooper I ended up utilizing another formula and I got the same results. Do you mind explaining me another question?

OpenStudy (issy14):

I came up with the following: w=theta/t and Vt= wr I set them equal to each other since they cover the same time, however, different distances. w=theta/t=wr then I cancelled out the w, and got t= theta/r, the problem is that my units don't add up .

OpenStudy (anonymous):

linear velocity= angular velocity * radius hence v= 188 * (35/2) cm/sec v = 3290 cm/sec = 32.9 m/sec